1650 students take an english exam. If 42% fail the exam, how many students failed the exam? how many students passed.
(could you show me how to do this problem?)

Answer by Earlsdon(6287)

To find 42% of 1650 you first convert the percentage to its decimal eqivalent:
42% = 42/100 = 0.42 then you multiply this by 1650.
0.42(1650) = 693
693 students failed the exam.
To find the number of students who passed, subtract the number who failed (693) from the total number of students (1650).
1650-693 = 957
957 students passed the exam.